# MICL: Improving In-Context Learning through Multiple-Label Words in Demonstration

### v1

@@ 4 Experiments
In this section, we examine the capacities of multiple label words in ICL from five perspectives: (1) 1-shot ICL classification performance with/without multiple label words insertion in baseline models and MICL (Sec. 4.2); (2) 5-shot ICL (Sec. 4.3); (3) The impact of leveraging extra label word mappings in prediction with/without it appearing in demonstrations (Sec. 4.4); (4) The effectiveness of MICL demonstration order compared to enumerating other permutations in ICL (Sec. 4.5); and (5) The influence of MICL with/without label-balanced demonstration (Sec. 4.6).
@@ 4.1 Setups
Datasets We evaluate our method on seven datasets, including SST-2 (Socher et al., 2013), CR (Ding et al., 2008), IMDB (Maas et al., 2011), ISEAR (Scherer and Wallbott, 1994), AMAN (Aman and Szpakowicz, 2008), TREC-6 (Li and Roth, 2002) and AGNews Zhang et al. (2015). We adopt the templates in Wang et al. (2023). The initial label word sets are from Hu et al. (2022) and Zhu et al. (2024). Detailed information on the datasets, templates and label sets is given in Appx. B.
@@
For the baseline models, including vanilla Llama2-7b, vanilla GPT2-xl, Topk Liu et al. (2022), SelfICL Wu et al. (2023), and DataICL Chang and Jia (2023), we adopt their proposed demonstration samples and paired them with class name and our multiple label words, respectively, to create sample-label pairs. Detailed information on the baseline models and their experiment settings is provided in Appx. C.
For each model, multiple-label word insertion is evaluated on validation set, which is the full training set excluding the samples selected for demonstrations. In Vanilla, TopK, and SelfICL, the full training set is split into a reduced training set and a validation set in an 8:2 ratio as demonstration samples vary with test samples. The demonstration samples are selected based on reduced training set, while the multiple-label word insertion is based on validation set. The ICL performance is evaluated on the same test set for all experiments. Results using class names in demonstrations are predicted based on the maximum logits over class names.
@@ 4.2 Main Results
+Demo-MLabels_CN 93.74$\uparrow$ 94.41$\uparrow$ 95.89$\uparrow$ 69.88$\uparrow$ 59.75$\uparrow$ 71.48$\uparrow$ 83.36$\uparrow$ 81.22
+Demo-MLabels_LW 93.74 94.41 95.89 72.88$\uparrow$ 59.05 71.03 83.88$\uparrow$ 81.55
@@
+Demo-MLabels_CN 93.52$\uparrow$ 93.55$\uparrow$ 94.69$\uparrow$ 83.20$\uparrow$ 58.15$\uparrow$ 67.38$\uparrow$ 79.59$\uparrow$ 81.44
+Demo-MLabels_LW 93.52 93.55 94.69 83.20 58.65$\uparrow$ 66.98 79.59 81.45
@@
+Demo-MLabels_CN 92.53$\uparrow$ 92.82$\uparrow$ 95.10$\uparrow$ 79.00$\uparrow$ 58.65$\uparrow$ 69.10$\uparrow$ 82.01$\uparrow$ 81.39
+Demo-MLabels_LW 92.53 92.82 95.10 82.80$\uparrow$ 58.65 69.30$\uparrow$ 82.01 81.96
@@
+Demo-MLabels_CN 95.28$\uparrow$ 92.55$\uparrow$ 94.70$\uparrow$ 74.40$\uparrow$ 55.64$\uparrow$ 71.30$\uparrow$ 84.83$\uparrow$ 81.40
+Demo-MLabels_LW 95.28 92.55 94.70 78.80$\uparrow$ 54.89 71.10 85.45$\uparrow$ 81.95
MICL 95.391 94.411 95.401 78.401 59.902 72.492 84.112 82.87
+Demo-MLabels_CN 95.971$\uparrow$ 95.151$\uparrow$ 95.601$\uparrow$ 79.801$\uparrow$ 65.162$\uparrow$ 73.552$\uparrow$ 86.552$\uparrow$ 84.54
+Demo-MLabels_LW 95.971 95.151 95.601 80.601$\uparrow$ 69.402$\uparrow$ 73.092 86.582$\uparrow$ 85.20
@@
+Demo-MLabels_CN 85.63$\uparrow$ 67.07$\uparrow$ 68.23$\uparrow$ 53.60$\uparrow$ 39.35$\uparrow$ 50.01$\uparrow$ 58.44$\uparrow$ 60.33
+Demo-MLabels_LW 85.63 67.07 70.13$\uparrow$ 54.76$\uparrow$ 39.30 50.56$\uparrow$ 58.24 60.81
@@
+Demo-MLabels_CN 84.51$\uparrow$ 66.22$\uparrow$ 65.47$\uparrow$ 60.80$\uparrow$ 40.85$\uparrow$ 54.75$\uparrow$ 55.24$\uparrow$ 61.12
+Demo-MLabels_LW 84.51 66.22 67.07$\uparrow$ 61.80$\uparrow$ 40.60 54.88$\uparrow$ 55.24 61.47
@@
+Demo-MLabels_CN 83.80$\uparrow$ 64.89 61.46$\uparrow$ 64.20$\uparrow$ 43.36$\uparrow$ 54.82$\uparrow$ 56.42$\uparrow$ 61.28
+Demo-MLabels_LW 83.80 64.89 62.16$\uparrow$ 65.40$\uparrow$ 43.11 55.08$\uparrow$ 56.42 61.55
@@
+Demo-MLabels_CN 84.84$\uparrow$ 63.83 69.80$\uparrow$ 58.20$\uparrow$ 36.84$\uparrow$ 48.04$\uparrow$ 51.14$\uparrow$ 58.96
+Demo-MLabels_LW 84.84 63.83 69.20 65.20$\uparrow$ 36.84 48.04 51.14 59.87
MICL 85.172 64.891 71.501 70.002 47.872 48.642 78.922 66.71
+Demo-MLabels_CN 91.652$\uparrow$ 65.961$\uparrow$ 73.401$\uparrow$ 70.402$\uparrow$ 49.622$\uparrow$ 59.732$\uparrow$ 79.082$\uparrow$ 70.55
+Demo-MLabels_LW 91.652 65.961 73.401 70.402 49.872$\uparrow$ 58.742 79.492$\uparrow$ 70.50
Table 1: ICL Experimental Results: ‘+Demo-MLabels_CN’ refers to the class-name result with multiple-label words enhanced in the demonstration (predicted based on the maximum logits over class names), and ‘+Demo-MLabels_LW’ refers to the label-words result with multiple-label words enhanced in the demonstration (predicted based on the maximum logits over inserted label words). The best accuracy results (%) are marked in bold. Marker 1 indicates results given under Eq 5, while marker 2 indicates results given under Eq 6. Upward arrow ($\uparrow$) in ‘+Demo-MLabels_CN’ signifies an increase in performance compared to the original method, while in ‘+Demo-MLabels_LW’, it signifies an increase in performance compared to ‘+Demo-MLabels_CN’.
We next present the results of each model and MICL (ours) under two settings: with and without multiple-label word insertion in 1-shot ICL. The label words in the initial sample-label pairs are updated label words in MICL and DataICL, while other models use class names. The best test accuracy achieved with multiple-label words inserted in demonstrations before meeting the stopping criterion (validation accuracy decreases with the insertion of label words) is reported as the result of ‘+Demo-MLabels’. Fig. 3 shows validation and test performance across different numbers of inserted label words (N) for each dataset in MICL. Details on inserted word settings in baseline models and ours for 1-shot settings are given in Appx. D.
@@
Figure 3: Validation and test performance under different label word quantity (N) in sample-multiple-label pairs. The red cross marks the reported result setting. In Llama2-7b, N is 2 for SST2, IMDB, TREC, ISEAR, and AGNews, 4 for CR, and 5 for AMAN. In GPT2-xl, N is 7 for ISEAR and 2 for others. The remaining datasets are in Appx. D.
The results in Table 1 lead to two conclusions:
1. Integrating multiple label words in demonstrations significantly enhances ICL performance. Across all baseline models, the use of multiple label words led to an average accuracy improvement of over 2% in Llama2-7b (with a maximum of 2.60% in SelfICL) and over 5% in GPT2-xl (with a maximum of 7.37% in vanilla-GPT2-xl).
2. MICL outperforms baselines across all datasets under the initial label word. It achieves an average accuracy improvement of 3.11% in Llama2-7b and 11.58% in GPT2-xl. Notably, the improvement in multi-class classification tasks is more significant: with accuracy gains of 6.02% (AMAN, Llama2-7b), 12.80% (TREC, GPT2-xl), 12.53% (AMAN, GPT2-xl), and 23.68% (AGNews, GPT2-xl).
By using multiple label words, MICL has an average accuracy improvement of 2.33% in Llama2-7b (with a maximum of 9.60% in AMAN) and 3.84% in GPT2-xl (with a maximum of 11.09% in ISEAR) compared with the initial setting where a single label word is used in sample-label pairing.
Enhancement via initial label words updates To ensure the quality of initial sample-label pairs, we update certain labels in two datasets based on the samples selected in MICL and DataICL before ICL experiments [^3] Vanilla, TopK, and SelfICL demonstration samples vary from test samples, leading to non-uniform sample-label pairs among the test samples in label evaluation. Consequently, we apply class names as the initial label words in the multiple-label word insertion experiment.. In AMAN, ‘other’ is replaced with ‘neutral’. In TREC, ‘entity’ is replaced with ‘animal’, ‘description’ with ‘definition’, ‘human’ with ‘persons’, ‘location’ with ‘state’, and ‘number’ with ‘numeric’. The enhanced accuracy presented in Table 2 demonstrates the superiority of our initial label word replacement method over the direct use of class names for improving 1-shot ICL.
@@
DataICL MICL DataICL MICL
@@
TREC 70.20 71.80 70.40 78.40 37.40 57.20 54.20 61.60
Table 2: Enhancement in ICL accuracy (%) through label evaluation and updates.
4.3 Effectiveness of MICL in 5-shot ICL
To explore the capability of multiple-label words further, we assess their effectiveness in a 5-shot setting using our method and SelfICL for one binary classification task (SST2) and one multi-class classification task (AMAN). For both methods, the 5-shot sample-label pairs are selected based on the top-5 scoring samples in the training set for each label. The validation set is created by removing all selected samples from the training set. The sample-label pairs are ordered based on the highest score achieved by each label among the selected samples.
SST2 AMAN Avg. SST2 AMAN Avg.
@@ 4.3 Effectiveness of MICL in 5-shot ICL
SelfICL 93.63 50.13 71.88 73.04 34.56 53.80
+Demo-MLabels_CN 94.89$\uparrow$ 55.89$\uparrow$ 75.39 86.60$\uparrow$ 40.35$\uparrow$ 63.48
+Demo-MLabels_LW 94.89 56.14$\uparrow$ 75.52 86.60 40.60$\uparrow$ 63.60
MICL 94.56 56.14 75.35 81.16 34.84 58.00
+Demo-MLabels_CN 95.39$\uparrow$ 58.40$\uparrow$ 76.90 89.18$\uparrow$ 51.80$\uparrow$ 70.49
+Demo-MLabels_LW 95.39 58.40 76.90 89.18 51.80 70.49
Table 3: Multiple-label word insertion accuracy (%) in 5-Shot ICL: Scoring method and test sets as in Table 1.
As shown in Table 3, the insertion of multiple label words is effective in 5-shot settings, yielding an average accuracy improvement of 3.64% and 9.80% in SelfICL, and 1.55% and 12.49% in MICL under Llama2-7b and GPT2-xl, respectively. MICL outperforms SelfICL across all results. Details on inserted words in baseline models and ours for 5-shot settings are in Appx. D, Table 11.
4.4 Analysis of Utilizing Multiple Label Words Mapping in Prediction
Table 1 and 3 show that ‘+Demo-MLabels_LW’ (prediction based on the maximum logit over inserted label words) achieves higher accuracy than ‘+Demo-MLabels_CN’ (prediction based on the maximum logit over class names) in some datasets. This seems to align with the idea that incorporating extra label word mapping in the final prediction can enhance prompt-based methods Schick and Schütze (2021). However, in ICL, models tend to predict based on the labels provided in demonstrations, which may perform differently with extra label word mappings in prediction. We compare the effectiveness of using extra label word mappings only in predictions versus including them in demonstrations for both binary-class and multi-class tasks. Additionally, we evaluate the impact of leveraging extra label word mappings under zero-shot settings.
SST2 TREC AMAN ISEAR AGNews Avg.
Llama2-7b
ZSL 88.96 68.80 48.87 58.60 67.29 66.50
ZSL_Sr 90.94 60.00$\downarrow$ 47.37$\downarrow$ 60.80 60.03$\downarrow$ 63.83
MICL 95.39 78.40 59.90 72.49 78.06 76.85
MICL_Sr 95.44 72.00$\downarrow$ 63.13 71.56$\downarrow$ 74.37$\downarrow$ 75.30
+Demo-MLabels_CN 95.97 79.80 65.16 73.55 86.55 80.21
+Demo-MLabels_LW 95.97 80.60 69.40 73.09 86.58 81.13
+Demo_MLabels_Sr 94.51$\downarrow$ 74.40$\downarrow$ 67.64$\downarrow$ 71.56$\downarrow$ 75.33$\downarrow$ 76.69
GPT2-xl
ZSL 79.57 38.00 39.60 42.33 53.16 50.53
ZSL_Sr 50.30$\downarrow$ 46.80 37.84$\downarrow$ 42.52 51.29$\downarrow$ 45.75
MICL_C 85.17 61.60 47.87 48.64 78.92 64.44
MICL_Sr 85.17 67.60 47.62$\downarrow$ 48.70 65.28$\downarrow$ 62.87
+Demo-MLabels_CN 91.65 70.40 49.62 59.73 79.08 70.10
+Demo-MLabels_LW 91.65 70.40 49.87 58.74 79.49 70.03
+Demo_MLabels_Sr 91.65 68.80$\downarrow$ 49.87 58.64$\downarrow$ 67.58$\downarrow$ 67.31
Table 4: Impact of prediction under class names, multiple-label words in demonstration and $S_{r}$. Arrow $\downarrow$ indicates a decrease in accuracy of predictions over $S_{r}$, compared to those over class names (CN) or inserted label words (LW).
In Table 4, most experiments show decreased performance when predicting based on label words in $S_{r}$ compared to predictions on label words that appeared in demonstrations (including class names). This suggests that simply applying extra label word mappings in the final prediction might disrupt information learned from demonstrations in ICL. Surprisingly, even under zero-shot learning (ZSL), where label information isn’t prompted, adding extra label knowledge still decreases some classification performance. These results highlight the complex role of labels in ICL classification.
4.5 Effectiveness of MICL Ordering
In MICL, we use LLM’s feedback over label words as a decision feature to order sample-label pairs. Although this approach may not provide the optimal ordering compared to evaluating all possible permutations, the computational cost of enumerating all possibilities in multi-class tasks is prohibitive. To assess the effectiveness of our ordering method, we compare the classification performance using MICL’s initial sample-label pairs against 30 and 50 random permutations (excluding MICL’s order) in multi-class tasks in Llama2-7b and GPT2-xl, respectively, and the flipped order in binary tasks.

### v4

@@ 4 Experiments
In this section, we examine the capacities of our method in ICL from the following perspectives: (1) 1-shot ICL classification comparison between baseline models and LICL (Sec.4.2); (2) 5-shot ICL (Sec.4.3); (3) Ablation studies on sample ordering (Sec.4.4.1) and multiple class-related words selection and ordering (Sec.4.4.2); (4) Scalability to larger LLMs (Sec.4.5), (5) General applicability of selected multiple class-related words (Sec.4.6); and (6) Further analyses (Sec.4.7).
@@ 4.1 Setups
Datasets We evaluate our method on seven datasets, including SST-2 (Socher et al., 2013), CR (Ding et al., 2008), IMDB (Maas et al., 2011), ISEAR (Scherer and Wallbott, 1994), AMAN (Aman and Szpakowicz, 2008), TREC-6 (Li and Roth, 2002) and AGNews Zhang et al. (2015). We adopt the templates in Wang et al. (2023). The initial class-related word pools are from Hu et al. (2022) and Zhu et al. (2024). Detailed information on the datasets, templates and label sets is given in Appx.D.
@@
For the baseline models, including vanilla LLaMA2-7b, vanilla GPT2-xl, Topk Liu et al. (2022), SelfICL Wu et al. (2023), and DataICL Chang and Jia (2023), we adopt their proposed demonstration samples and pair them with class name and our multiple class-related words, respectively, to create sample-label pairs. Detailed descriptions of the baseline models, including comparing their methods with ours and an overview of their experimental settings, are in Appx.E.
For each model, multiple class-related words insertion is evaluated on validation set, which is the full training set excluding the samples selected for demonstrations. In Vanilla, TopK, and SelfICL, the full training set is split into a reduced training set and a validation set in an 8:2 ratio as demonstration samples vary with test samples. The demonstration samples are selected based on the reduced training set, while the multiple class-related words insertion is based on validation set. The ICL performance is evaluated on the same test set for all experiments.
@@ 4.2 Main Results
We present the results of each model and LICL (ours) under two settings: with and without multiple class-related words insertion in 1-shot ICL.
@@
+MLabels_CN 93.74$\uparrow$ 94.41$\uparrow$ 95.89$\uparrow$ 69.88$\uparrow$ 59.75$\uparrow$ 71.48$\uparrow$ 83.36$\uparrow$ 81.22
+MLabels_LW 93.74 94.41 95.89 72.88$\uparrow$ 59.05 71.03 83.88$\uparrow$ 81.55
@@
+MLabels_CN 93.52$\uparrow$ 93.55$\uparrow$ 94.69$\uparrow$ 83.20$\uparrow$ 58.15$\uparrow$ 67.38$\uparrow$ 79.59$\uparrow$ 81.44
+MLabels_LW 93.52 93.55 94.69 83.20 58.65$\uparrow$ 66.98 79.59 81.45
@@
+MLabels_CN 92.53$\uparrow$ 92.82$\uparrow$ 95.10$\uparrow$ 79.00$\uparrow$ 58.65$\uparrow$ 69.10$\uparrow$ 82.01$\uparrow$ 81.39
+MLabels_LW 92.53 92.82 95.10 82.80$\uparrow$ 58.65 69.30$\uparrow$ 82.01 81.96
@@
+MLabels_CN 95.28$\uparrow$ 92.55$\uparrow$ 94.70$\uparrow$ 74.40$\uparrow$ 55.64$\uparrow$ 71.30$\uparrow$ 84.83$\uparrow$ 81.40
+MLabels_LW 95.28 92.55 94.70 78.80$\uparrow$ 54.89 71.10 85.45$\uparrow$ 81.95
LICL 95.391 94.411 95.401 78.401 59.902 72.492 84.112 82.87
+MLabels_CN 95.971$\uparrow$ 95.151$\uparrow$ 95.601$\uparrow$ 79.801$\uparrow$ 65.162$\uparrow$ 73.552$\uparrow$ 86.552$\uparrow$ 84.54
+MLabels_LW 95.971 95.151 95.601 80.601$\uparrow$ 69.402$\uparrow$ 73.092 86.582$\uparrow$ 85.20
@@
+MLabels_CN 85.63$\uparrow$ 67.07$\uparrow$ 68.23$\uparrow$ 53.60$\uparrow$ 39.35$\uparrow$ 50.01$\uparrow$ 58.44$\uparrow$ 60.40
+MLabels_LW 85.63 67.07 70.13$\uparrow$ 54.76$\uparrow$ 39.30 50.56$\uparrow$ 58.24 60.81
@@
+MLabels_CN 84.51$\uparrow$ 66.22$\uparrow$ 65.47$\uparrow$ 60.80$\uparrow$ 40.85$\uparrow$ 54.75$\uparrow$ 55.24$\uparrow$ 61.12
+MLabels_LW 84.51 66.22 67.07$\uparrow$ 61.80$\uparrow$ 40.60 54.88$\uparrow$ 55.24 61.47
@@
+MLabels_CN 83.80$\uparrow$ 64.89 61.46$\uparrow$ 64.20$\uparrow$ 43.36$\uparrow$ 54.82$\uparrow$ 56.42$\uparrow$ 61.28
+MLabels_LW 83.80 64.89 62.16$\uparrow$ 65.40$\uparrow$ 43.11 55.08$\uparrow$ 56.42 61.55
@@
+MLabels_CN 84.84$\uparrow$ 63.83 69.80$\uparrow$ 58.20$\uparrow$ 36.84$\uparrow$ 48.04$\uparrow$ 51.14$\uparrow$ 58.96
+MLabels_LW 84.84 63.83 69.20 65.20$\uparrow$ 36.84 48.04 51.14 59.87
LICL 85.172 64.891 71.501 70.002 47.872 48.642 78.922 66.71
+MLabels_CN 91.652$\uparrow$ 69.951$\uparrow$ 73.401$\uparrow$ 70.402$\uparrow$ 49.622$\uparrow$ 59.732$\uparrow$ 79.082$\uparrow$ 70.55
+MLabels_LW 91.652 69.951 73.401 70.402 49.872$\uparrow$ 58.742 79.492$\uparrow$ 70.50
Table 1: ICL Experimental Results: LICL is demonstrated on initial sample-label pairs while ‘+MLabels’ is on sample-multiple-label pairs. ‘+MLabels_CN’ refers to prediction based on the maximum logit over class names, and ‘+MLabels_LW’ refers to prediction based on the maximum logit over inserted class-related words. The best accuracy results (%) are marked in bold. Marker 1 indicates results given under Eq.1, while marker 2 indicates results given under Eq.2. Upward arrow ($\uparrow$) in ‘+MLabels_CN’ signifies an increase in performance compared to the original method, while in ‘+MLabels_LW’, it signifies an increase in performance compared to ‘+MLabels_CN’.
LICL Outperforms Over Baselines: As demonstrated in Table 1, the proposed LICL consistently outperforms the baseline methods across all datasets when evaluated under initial class-related word conditions. It achieves an average accuracy improvement of 3.11% in LLaMA2-7b and 11.58% in GPT2-xl. Particularly notable are the gains in multi-class classification tasks, where it records accuracy improvements of 6.02% (AMAN, LLaMA2-7b), 12.8% (TREC, GPT2-xl), 12.53% (AMAN, GPT2-xl), and 23.68% (AGNews, GPT2-xl).
Multiple Class-Related Words Enhanced: Utilizing multiple class-related words (‘+MLabels’) significantly improves performance across all models. This strategy results in an average baselines’ accuracy increase of over 2% in LLaMA2-7b, reaching up to 2.6% in SelfICL, and more than 5% in GPT2-xl, with a maximum improvement of 7.37% in vanilla GPT2-xl. Compared to configurations that employ a single class-related word per sample-label pair, LICL achieves notable accuracy gains, with increases of 2.33% in LLaMA2-7b (peaking at 9.6% in AMAN) and 3.84% in GPT2-xl (peaking at 11.09% in ISEAR). These improvements demonstrate significant enhancements over the initial single-label setup, underscoring the efficacy of the multiple class-related word approach.
Validation-Testing Results Alignment: Fig.3 displays the validation and test performance across various counts of inserted class-related words (N) for each dataset in LICL. Generally, the trends in test performance are consistent with those in validation performance. The peak test accuracy was achieved by incorporating multiple class-related words before reaching the stopping criterion, at which point the validation accuracy begins to decrease with further insertion of class-related words. This consistency supports the reliability of our stopping criterion.
@@
Figure 3: Validation and test performance under inserted word quantity (N) in sample-multiple-label pairs. The red cross marks the reported result setting. In LLaMA2-7b, N is 2 for SST2, IMDB, TREC, ISEAR, and AGNews, 4 for CR, and 5 for AMAN. In GPT2-xl, N is 7 for ISEAR and 2 for others. The remaining datasets are in Appx.F.
Initial Class-Related Word Updates Improved To improve the quality of the initial sample-label pairs, we jointly update certain labels in two datasets based on the samples selected in LICL and DataICL before ICL experiments [^4] Updating initial class-related words ensures they have a high logit value relative to other words with the same label polarity to provide effective class-related guidance in demonstrations for all testing samples. However, Vanilla, TopK, and SelfICL use varying demonstration samples across test instances, leading to inconsistent evaluations and making updates impractical. Therefore, we use class names as initial words in the multiple insertion experiment.. In AMAN, ‘other’ is replaced with ‘neutral’. In TREC, ‘entity’ is replaced with ‘animal’, ‘description’ with ‘definition’, ‘human’ with ‘persons’, ‘location’ with ‘state’, and ‘number’ with ‘numeric’. The enhanced accuracy presented in Table 2 highlights the effectiveness of class-related word updates over the direct use of predefined class names.
@@
DataICL LICL DataICL LICL
@@
TREC 70.20 71.80 70.40 78.40 37.40 57.20 54.20 70.00
Table 2: Enhancement in 1-shot ICL accuracy (%) through label evaluation and updates.
4.3 Effectiveness of LICL in 5-shot ICL
To explore the capability of multiple class-related words further, we assess their effectiveness in a 5-shot setting using our method and SelfICL for one binary classification task (SST2) and one multi-class classification task (AMAN). For both methods, the 5-shot sample-label pairs are selected based on the top-5 scoring samples in the training set for each label. The validation set is created by removing all selected samples from the training set. The sample-label pairs are ordered based on the highest score achieved by each label among the selected samples.
SST2 TREC AMAN Avg. SST2 TREC AMAN Avg.
@@ 4.3 Effectiveness of LICL in 5-shot ICL
SelfICL 93.63 73.20 50.13 72.32 73.04 59.00 34.56 55.53
+MLabels_CN 94.89 74.00 55.89 74.93 86.60 69.20 40.4 65.38
+MLabels_LW 94.89 74.00 56.14 75.01 86.60 69.20 40.60 65.47
LICL 94.56 76.20 56.14 75.35 81.16 72.00 34.84 62.67
+MLabels_CN 95.39 78.60 58.40 76.90 89.18 74.40 51.80 71.79
+MLabels_LW 95.39 78.60 58.40 76.90 89.18 73.00 51.80 71.33
Table 3: Multiple class-related words insertion accuracy (%) in 5-shot ICL: Scoring method and test sets as in Table 1.
As shown in Table 3, the insertion of multiple class-related words is effective in 5-shot settings, yielding an average accuracy improvement of 2.69% and 9.93% in SelfICL, and 1.08% and 9.13% in LICL under LLaMA2-7b and GPT2-xl. LICL outperforms SelfICL across all results. Inserted words for both baseline and our models are in Appx.F, and enhancements for updating initial class-related words under 5-shot settings are in Appx.G.1.
4.4 Ablation Studies
4.4.1 Effectiveness of LICL Ordering
In LICL, we use LLM’s feedback over class-related words as a decision feature to order sample-label pairs. To evaluate our ordering method’s effectiveness, we compare 1-shot ICL classification performance using LICL’s initial sample-label pairs against 30 (LLaMA2-7b) and 50 (GPT2-xl) random permutations (excluding LICL’s order) in multi-class tasks [^5] We randomize the order of samples selected by LICL to create various permutations and then randomly select 30 or 50 different permutations from the generated pool. Each permutation is assessed individually in a 1-shot ICL classification., and a flipped order in binary tasks.